### v1

# Technology prediction of a 3D model using Neural Network

Abstract
Accurate estimation of production times is critical for effective manufacturing scheduling, yet traditional methods relying on expert analysis or historical data often fall short in dynamic or customized production environments. This paper introduces a data-driven approach that predicts manufacturing steps and their durations directly from a product’s 3D model. By rendering the model into multiple 2D images and leveraging a neural network inspired by the Generative Query Network, the method learns to map geometric features into time estimates for predefined production steps enabling scalable, adaptive, and precise process planning across varied product types.
keywords:
Technology Prediction , Neural Network , Machine Learning , Deep Learning
journal:
[…]
[m]organization=Masta Solutions sp. z o.o.,city=Poznan, country=Poland
[…]
[p]organization=Institute of Computing Science, Poznan University of Technology,city=Poznan, country=Poland
1 Introduction
Accurate production scheduling is a cornerstone of efficient manufacturing. In practice, schedules are generated based on estimates of processing times required for each step in the production process. However, when these estimates deviate from actual conditions—due to missing or outdated data - the generated schedules quickly become obsolete. This mismatch leads to inefficiencies, frequent rescheduling, and, ultimately, a decline in overall manufacturing performance.[1]
Traditionally, estimating production times has required manual analysis by experts or historical data from similar products. While effective in some cases, these methods are time-consuming, error-prone, and difficult to scale, particularly in environments where products are frequently customized or modified. For instance, Żywicki and Osiński discuss the challenges manufacturers face in calculating production times for customized products, highlighting the lack of real data on standard times required for individual manufacturing operations. [2, 3]
This research proposes an automated approach to predicting production steps and their corresponding times based solely on the 3D model of a product. The central idea is to train a neural network that learns to map geometric features of a product - extracted from rendered images of its 3D model - into a vector of processing times across a predefined set of manufacturing steps.
The proposed method draws inspiration from the Generative Query Network (GQN)[4, 5], representing 3D models as collections of 2D images taken from various viewpoints. These images serve as input to an encoder-decoder neural network architecture: the encoder, based on EfficientNetB0 architecture [6], extracts meaningful latent representations from individual images, and the decoder aggregates and transforms these into time estimates for each production step. This architecture is flexible, allowing the use of a variable number of input images, and scalable to different product types without manual feature engineering.
By learning directly from historical manufacturing data, this model offers a promising direction toward data-driven process planning, enabling faster, more adaptive, and more precise scheduling in modern production systems.
2 Methods
2.1 Task
The aim is to, based on a 3D model ($m_{i}$) of an item $i$ predict a vector $t_{i}$ of length $k$ denoting the times required at each step $j,j\in\{0,1,...,k-1\}$ to produce the item. The $k$ steps are predefined and constant for each $i$. The items can be divided into a reference set $A^{R}$ for which both $m$ and $t$ are available and a non-reference set $A^{NR}$ for which only the 3D model is available but times are unknown. The reference set consists of items already produced or analyzed by experts, hence the production times and steps are known. Formally, the task is to create a function $f$ that will map $m$ into $t$, $f(m)=t$.
2.2 Data Processing
The proposed architecture is inspired by the Generative Query Network (GQN) [4] approach. In that model, a 3D scene is represented as a set of classical 2D images picturing the scene from different angles. That procedure was adopted for the technology prediction task. The process is presented in Figure LABEL:fig:images. To represent an item to be produced a set of images $I=\{i_{0},i_{1},\ldots\}$ is generated. They are taken from different camera positions $P=\{p_{0},p_{1},\ldots\}$ to capture each detail of the item. This is possible due to the characteristics of the considered objects - all components visible from the outside. In the presented approach the number of images was randomly selected between $3$ and $5$. To describe the position of a camera $3$ numbers were used denoting it’s position alongside $x$, $y$, and $z$ axes and additional $3$ numbers informing about it’s rotation on these axes.
2.3 Neural architecture
The neural architecture of the proposed model, presented in Figure LABEL:fig:neural, is based on the encoder-decoder approach. The former is responsible for transforming an image into a latent vector [7] compressing relevant data from the image to a fixed-size series of numbers. The latter is used to change the latent vectors into meaningful predictions, in this case, times of production. In the proposed model an EfficientNetB0 [6], a state-of-the-art machine learning model for image processing tasks, neural network is used as the encoder while a simple multilayer perceptron network with 3 layers serves as the decoder. The encoder $f^{\prime}$ is applied independently to each input image, producing a latent vector for each of them $L=\{l_{0},l_{1},...\}$, $f^{\prime}(i_{i})=l_{i}$. Those vectors are then aggregated using a piecewise mean operator to create a vector $l$ containing information regarding the whole 3D model. This procedure allows to use any number of pictures within the same network since the architecture and the number of trainable parameters do not depend on the number of input images. This vector is then processed by the decoder $g$ consisting of simple classic fully connected layers. The last layer consists of $k$ neurons with a relu activation function turning the input into a predicted vector storing times required at particular production steps with time 00 denoting that a given step is not needed.
3 Experiment
The dataset in the experiment consists of 539 3D models of items with corresponding vectors denoting times of $k=6$ different manufacturing stages (welding, cutting, bending, screwing, drilling, and assembling). If producing a given item does not require all steps then for the unnecessary stages the time is equal to zero. Sample times are presented in Table 1 and selected objects are presented on Figure 11. $60$ of these objects were put aside as a test set used to evaluate the model, while the remaining $479$ items were used for training purposes. For each training object a hundred snapshots were generated. During the training in each epoch a different subset was selected to represent an item. This procedure was crucial due to a very low number of training examples. In this data augmentation process number of unique examples coming from one item was close to 80 million (${100\choose 3}+{100\choose 4}+{100\choose 5}=79370445$. However, they cannot be treated as independent training examples since they all originate from the same 3D object, nevertheless, such procedure allows preventing overfitting of the model, thus attaining better results.
Table 1: Example matrix presenting manufacturing times of selected items
ID welding cutting bending screwing drilling assembling
1 0 0 20 0 200 0
2 300 60 0 0 0 0
3 0 30 5 0 15 0
4 0 60 0 0 0 60
5 0 0 10 0 260 0
6 60 30 0 0 30 0
7 0 30 0 60 0 0
8 300 60 0 0 0 0
9 0 60 0 0 0 300
1 2 3
Figure 11: Example snapshots of objects used to train the model.
Table 2: Performance Comparison of Two Models
Model MAE MAPE
GQN 2.76 7.8
PointNet 16 18.66
To verify the quality of the proposed approach a PointNet [8] model was trained on the point clouds generated from the 3D models of the same dataset. A similar approach was successfully applied to such problems as detecting geometric defects in gears [9], identification of mechanical parts for robotic disassembly [10], or classification of crystalline structures in complex plasmas [11]. For this model and data type, there was no straightforward way to augment the data, hence the PointNet model training was significantly faster but the model was not able to learn how to map a point cloud to production times. Results obtained by both models are presented in Table 2. The Mean Absolute Error (MAE) for results obtained using the GQN approach is equal to 2.76 seconds which is less than a resolution of data since all times were rounded to 5 seconds. The Mean Absolute Percentage Error (MAPE) of this model is equal to 7.6 percentage points. To calculate MAPE when the expected value is zero the following procedure was applied
$mape(real,pred)=\begin{cases}0,&\text{if }real=0\land pred<1\\ 1,&\text{if }real=0\land pred\geq 1\\ \frac{|real-pred|}{real},&\text{otherwise}\end{cases}$
meaning predicting more than 1 second if the real value is zero results in 100% error while predicting less than a second results in 0% error. The pointnet model was over 5 times worse regarding MAE and over 2 times worse regarding MAPE.
4 Conclusions
The proposed neural network is capable of learning the required steps and their durations to produce an item and outperforms the PointNet model. Such results were possible due to the characteristics of the considered items i.e. simple shapes with no hidden details. In the case of more complex objects, the approach based on representing an item using some snapshot would not allow to capture all necessary information. The second crucial aspect was the limited number of training examples. Regarding this aspect, the proposed approach with a straightforward data augmentation strategy had a huge advantage. Further study with significantly more 3D models including those more complex and with hidden details is planned. In that scenario, more complex approaches will be considered e.g. extending the current approach by using not only snapshots but also cross sections and comparing it with a PointNet++ [12] model.
CRediT authorship contribution statement
Grzegorz Miebs: Writing - Original Draft, Investigation, Methodology, Software. Rafał  Bachorz: Writing - Review & Editing, Investigation, Methodology, Software.
Declaration of competing interest
The authors declare that they have no known competing financial interests or personal relationships that could have appeared to influence the work reported in this paper.
Acknowledgements
We thank Patrycja Nowak for designing the graphics used in the paper. This project has received funding from the European Union’s Horizon 2020 research and innovation programme under grant agreement N. 101017057